ord -Exam2 FallzuT4verzaprn.docx SHOW WORK TO GET FULL CREDIT Problem 7) Suppose that a certain random variable, X, has the fPlease show how to solve this.
Answer #1

i) P(X > 1) = 1 - P(X < 1) = 1 - F(1) = 1 - 0.7 = 0.3

ii) P(4 < X < 6) = F(4) - F(1) = 0.9 - 0.7 = 0.2